# Stabilizer Ranks, Barnes Wall Lattices and Magic Monotones

### v1

@@ 3 Linear Lower Bound on Stabilizer Rank
Let $\mathinner{|{\phi}\rangle}$ be a pure state such that $\mathinner{|{\phi}\rangle}=\sum_{j=1}^{k}c_{i}\mathinner{|{s_{j}}\rangle}$, where $\mathinner{|{s_{j}}\rangle}$’s are all linearly independent stabilizer states, $c_{j}\in\mathbb{C}$ with $\left\|\mathbf{c}\right\|_{1}\geq\alpha$ for some $\alpha>0$. Then we have that $\alpha\leq\sqrt{e}\cdot 2k^{\frac{k+1}{2}}$.
@@
We also know that the $\mathbb{Z}$-lattice generated by $\{\mathinner{|{s_{j}}\rangle}\}$ and $\{i\mathinner{|{s_{j}}\rangle}\}$ has minimal vectors of size $1$. Let $G$ be the Gram matrix for these $2k$ vectors (here, we consider the space $\mathbb{C}^{d}$ these vectors lie in as $\mathbb{R}^{2d}$, and this change does not alter length’s of vectors). By Minkowski’s Theorem (Theorem 28), we have that $\mathrm{det}(G)\geq\sqrt{2k}^{-2k}$.
@@
$(2k)^{-k}\leq\frac{2ek}{\alpha^{2}}\implies\alpha\leq\sqrt{e}\cdot 2k^{\frac{k+1}{2}}$
@@
Let $\mathinner{|{\phi}\rangle}$ be a pure state with stabilizer fidelity $F$. Suppose there is a state $\mathinner{|{\psi}\rangle}$ with stabilizer rank $k$ such that $\frac{|\mathinner{\langle{\phi|\psi}\rangle}|}{\sqrt{F}}\geq\beta$. Then we have that $\beta\leq\sqrt{e}(2k)^{\frac{k+1}{2}}$.
@@
Let $\mathinner{|{\psi}\rangle}$ have an exact stabilizer decomposition with $k$ linearly independent stabilizer states $\mathinner{|{s_{1}}\rangle}\dots\mathinner{|{s_{k}}\rangle}$. Let $V$ be the space spanned by these states, and let $P$ be the orthogonal projector onto this space. Without loss of generality we can assume $\mathinner{|{\psi}\rangle}=\frac{P\mathinner{|{\phi}\rangle}}{\left\|P\mathinner{|{\phi}\rangle}\right\|_{2}}$, as it is the vector with the largest inner product (in magnitude) with $\mathinner{|{\phi}\rangle}$ in $V$. With this, we have that $\left\|P\mathinner{|{\phi}\rangle}\right\|_{2}=|\mathinner{\langle{\psi|\phi}\rangle}|$. Now, we know that ${\rm{Tr~{}}}[{{\mathinner{|{\phi}\rangle}}{\mathinner{\langle{\phi}|}}}{{\mathinner{|{s_{j}}\rangle}}{\mathinner{\langle{s_{j}}|}}}]=|\mathinner{\langle{\phi|s_{j}}\rangle}|^{2}\leq F$ for each $j$. Therefore,
$|\mathinner{\langle{\psi|s_{j}}\rangle}|^{2}$ $={\rm{Tr~{}}}[{{\mathinner{|{\psi}\rangle}}{\mathinner{\langle{\psi}|}}}{{\mathinner{|{s_{j}}\rangle}}{\mathinner{\langle{s_{j}}|}}}]$ (3.2)
$={\rm{Tr~{}}}[(P{{\mathinner{|{\psi}\rangle}}{\mathinner{\langle{\psi}|}}}P){{\mathinner{|{s_{j}}\rangle}}{\mathinner{\langle{s_{j}}|}}}]\frac{1}{\left\|P\mathinner{|{\phi}\rangle}\right\|_{2}^{2}}$ (3.3)
$={\rm{Tr~{}}}[{{\mathinner{|{\psi}\rangle}}{\mathinner{\langle{\psi}|}}}(P{{\mathinner{|{s_{j}}\rangle}}{\mathinner{\langle{s_{j}}|}}}P)]\frac{1}{|\mathinner{\langle{\psi|\phi}\rangle}|^{2}}$ (3.4)
$={\rm{Tr~{}}}[{{\mathinner{|{\psi}\rangle}}{\mathinner{\langle{\psi}|}}}{{\mathinner{|{s_{j}}\rangle}}{\mathinner{\langle{s_{j}}|}}})]\frac{1}{|\mathinner{\langle{\psi|\phi}\rangle}|^{2}}$ (3.5)
@@
By Theorem 1, we get that $\beta\leq\sqrt{e}\cdot(2k)^{\frac{k+1}{2}}$. Now, if $\beta=2^{\Omega(n)}$, then $(k+1)\log(2k)=\Omega(n)\implies k=\Omega\left(\frac{n}{\log n}\right)$
@@ 4 Magic Monotones and Barnes Wall Lattices
$\mathcal{N}(\mathinner{|{\phi}\rangle})=\min\{||c\mathinner{|{\phi}\rangle}||_{2}^{2}:c\mathinner{|{\phi}\rangle}\in BW_{n},~{}c\in\mathbb{C^{*}}\}$
@@
Note that the $2$-norm is equivalent to the norm specified in [12] on $\mathbb{Q}[i]$. In our results we will restrict to $\mathbb{Q}[i]$.
@@ 4.1 Properties of $\mathcal{N}$ and $\mathcal{N}_{\delta}$
A vector $\mathinner{|{\phi}\rangle}$ is proportionally minimal in $BW_{n}$ if $\mathinner{|{\phi}\rangle}\in BW_{n}$ and for all $10<|c|<$ 1, $c\mathinner{|{\phi}\rangle}\notin BW_{n}$.
@@
$\forall a,b\in\mathbb{Z}[i]:ac\mathinner{|{\phi}\rangle}+b\\ ket{\phi}\in BW_{n}$
@@
    $\mathcal{N}$ and $\mathcal{N}_{\delta}$ is invariant under the action of the Clifford group (Also see [12]).
@@
(i) and (ii) follow directly from the invariance of $BW_{n}$ under the action of the Clifford group. [12].
@@
Now, (iv) follows directly from $(ii)$ and $(iii)$. For (v), we use a technique from [12]. For any state $\mathinner{|{\phi}\rangle}$, let $\mathinner{|{\phi^{\prime}}\rangle}$ be its proportionally minimal vector. Now, we can write it as, for some vectors $\mathinner{|{v}\rangle},\mathinner{|{w}\rangle}$,
@@
For each such $\alpha$, notice that $(1+i)\mathinner{|{0}\rangle}\mathinner{|{\alpha}\rangle}$ and $(1+i)\mathinner{|{1}\rangle}\mathinner{|{\alpha}\rangle}$ both lie in the dual of $BW_{n}$. Taking their inner products with $\mathinner{|{\phi^{\prime}}\rangle}$, we get that $-i\mathinner{\langle{\alpha|v}\rangle}$ and $-i\mathinner{\langle{\alpha|w}\rangle}$ respectively, proving that $\mathinner{|{v}\rangle},\mathinner{|{w}\rangle}\in BW_{n-1}$.
@@
Note that a similar property to Theorem 37 $(v)$ holds for the dyadic monotone [18].
@@ 4.2 Connection Between $\mathcal{N}_{\delta}$ and Stabilizer Ranks
$\forall\mathinner{|{\phi}\rangle}\in V~{}\exists~{}\mathinner{|{\psi}\rangle}\in L~{}\text{s.t}~{}||\mathinner{|{\phi}\rangle}-\mathinner{|{\psi}\rangle}||_{2}\leq\sqrt{k}$
@@
[…]
@@
$\forall\mathinner{|{\phi}\rangle}\in V~{}\exists~{}\mathinner{|{\psi}\rangle}\in L~{}\text{s.t}~{}||\mathinner{|{\phi}\rangle}-\mathinner{|{\psi}\rangle}||_{2}\leq\sqrt{2k}$
@@ 4.3 Relating $\mathcal{N}$ and $CS$-count for exact state preparation
Figure 1: $CS$-gate injection using the magic state $\mathinner{|{CS}\rangle}$.
@@
$\mathcal{N}(\mathinner{|{\psi}\rangle})|\mathcal{N}(\mathinner{|{CS}\rangle})^{m}=2^{m}$
@@ 5 Fidelity Amplification
$\frac{{\rm{Tr~{}}}[(I-{{\mathinner{|{\psi}\rangle}}{\mathinner{\langle{\psi}|}}})\,{{\mathinner{|{\psi^{\prime}}\rangle}}{\mathinner{\langle{\psi^{\prime}}|}}}]}{{\rm{Tr~{}}}[{{\mathinner{|{\psi}\rangle}}{\mathinner{\langle{\psi}|}}}\,{{\mathinner{|{\psi^{\prime}}\rangle}}{\mathinner{\langle{\psi^{\prime}}|}}}]}$ (5.1)
This can very easily be related to fidelity when $\left\|\mathinner{|{\phi^{\prime}}\rangle}\right\|_{2}=1$, and the expression turns out to be $\frac{1-F(\mathinner{|{\psi}\rangle},\mathinner{|{\phi}\rangle})}{F(\mathinner{|{\psi}\rangle},\mathinner{|{\phi}\rangle})}$
Observation 43.
@@
$\frac{{\rm{Tr~{}}}[(I-({{\mathinner{|{H}\rangle}}{\mathinner{\langle{H}|}}})^{\otimes n})\,{{\mathinner{|{\psi}\rangle}}{\mathinner{\langle{\psi}|}}}]}{{\rm{Tr~{}}}[({{\mathinner{|{H}\rangle}}{\mathinner{\langle{H}|}}})^{\otimes n}\,{{\mathinner{|{\psi}\rangle}}{\mathinner{\langle{\psi}|}}}]}$ (5.2)
@@
with ${\rm{Tr~{}}}[{{\mathinner{|{\phi}\rangle}}{\mathinner{\langle{\phi}|}}}]=\epsilon$, and $\mathinner{\langle{\phi}|}\mathinner{|{H}\rangle}^{\otimes n}=0$.
@@
with $\frac{I+C}{2}\mathinner{|{\phi}\rangle}$ orthogonal to $\mathinner{|{H}\rangle}^{\otimes n}$. Define the $LHS$ above to be $\mathinner{|{\psi^{\prime\prime}}\rangle}$. Note that $\chi(\mathinner{|{\psi^{\prime\prime}}\rangle})\leq 2\chi(\mathinner{|{\psi^{\prime}}\rangle})\leq 2k$. Also, note that ${\rm{Tr~{}}}[{{\mathinner{|{\psi^{\prime}}\rangle}}{\mathinner{\langle{\psi^{\prime}}|}}}\,({{\mathinner{|{H}\rangle}}{\mathinner{\langle{H}|}}})^{\otimes n}]={\rm{Tr~{}}}[{{\mathinner{|{\psi^{\prime\prime}}\rangle}}{\mathinner{\langle{\psi^{\prime\prime}}|}}}\,({{\mathinner{|{H}\rangle}}{\mathinner{\langle{H}|}}})^{\otimes n}]$. Thus, the denominator of Equation 5.2 is the same for both $\mathinner{|{\psi^{\prime}}\rangle}$ and $\mathinner{|{\psi^{\prime\prime}}\rangle}$.
Now, what we expect to happen is that $\frac{I+C}{2}\mathinner{|{\phi}\rangle}$ will have a lot of cancellations if written out in the $\mathinner{|{H}\rangle},\mathinner{|{H^{\perp}}\rangle}$ basis. This is because, for each such basis element, except $\mathinner{|{H}\rangle}^{\otimes n}$ (which by orthogonality has coefficient 00 in the expansion of $\mathinner{|{\phi}\rangle}$), it is an eigenvector of $\frac{I+C}{2}$, with the corresponding eigenvalue being $1$ or 00 with equal probability. We also know that for any vector $\mathinner{|{x}\rangle}$, we have that $\mathinner{\langle{x|x}\rangle}$ is the sum of squares of the coefficients of these basis elements. Thus, we have that for $\mathinner{|{\phi^{\prime}}\rangle}=\frac{I+C}{2}\mathinner{|{\phi}\rangle}$, $\mathbb{E}[\mathinner{\langle{\phi^{\prime}|\phi^{\prime}}\rangle}]=\frac{1}{2}\mathinner{\langle{\phi|\phi}\rangle}=\frac{\epsilon}{2}$, where the randomness is over the choices of $C$. Therefore, there exists a $C$ for which $\mathinner{\langle{\phi^{\prime}|\phi^{\prime}}\rangle}$ is at most $\frac{\epsilon}{2}$. This gives us our result. ∎
@@
Note that if we apply the above result to $\mathinner{|{0}\rangle}^{\otimes n}$, we retrieve back the $O(2^{0.23n})$ rank approximate stabilizer decomposition of $\mathinner{|{H}\rangle}^{\otimes n}$ from [6]. In fact, our technique can be thought of as generalization for their construction.

### v2

@@ 3 Linear Lower Bound on Stabilizer Rank
Let $\mathinner{|{\phi}\rangle}$ be a pure state such that $\mathinner{|{\phi}\rangle}=\sum_{j=1}^{k}c_{i}\mathinner{|{s_{j}}\rangle}$, where $\mathinner{|{s_{j}}\rangle}$’s are all linearly independent stabilizer states, $c_{j}\in\mathbb{C}$ with $\left\|\mathbf{c}\right\|_{1}\geq\alpha$ for some $\alpha>0$. Then we have that $\alpha\leq\sqrt{e}\cdot(2k)^{\frac{2k+1}{2}}$.
@@
We also know that the $\mathbb{Z}$-lattice generated by $\{\mathinner{|{s_{j}}\rangle}\}$ and $\{i\mathinner{|{s_{j}}\rangle}\}$ has minimal vectors of size $1$. Let $G$ be the Gram matrix for these $2k$ vectors (here, we consider the space $\mathbb{C}^{d}$ these vectors lie in as $\mathbb{R}^{2d}$, and this change does not alter length’s of vectors). By Minkowski’s Theorem (Theorem 30), we have that $\mathrm{det}(G)\geq({2k})^{-2k}$.
@@
$(2k)^{-2k}\leq\frac{2ek}{\alpha^{2}}\implies\alpha\leq\sqrt{e}\cdot(2k)^{\frac{2k+1}{2}}$
@@
Let $\mathinner{|{\phi}\rangle}$ be a pure state with stabilizer fidelity $F$. Suppose there is a state $\mathinner{|{\psi}\rangle}$ with stabilizer rank $k$ such that $\frac{|\mathinner{\langle{\phi|\psi}\rangle}|}{\sqrt{F}}\geq\beta$. Then we have that $\beta\leq\sqrt{e}(2k)^{\frac{2k+1}{2}}$.
@@
Let $\mathinner{|{\psi}\rangle}$ have an exact stabilizer decomposition with $k$ linearly independent stabilizer states $\mathinner{|{s_{1}}\rangle}\dots\mathinner{|{s_{k}}\rangle}$. Let $V$ be the space spanned by these states, and let $P$ be the orthogonal projector onto this space. Without loss of generality we can assume $\mathinner{|{\psi}\rangle}=\frac{P\mathinner{|{\phi}\rangle}}{\left\|P\mathinner{|{\phi}\rangle}\right\|_{2}}$, as it is the vector with the largest inner product (in magnitude) with $\mathinner{|{\phi}\rangle}$ in $V$. With this, we have that $\left\|P\mathinner{|{\phi}\rangle}\right\|_{2}=|\mathinner{\langle{\psi|\phi}\rangle}|$. Now, we know that ${\rm{Tr~}}[{{\mathinner{|{\phi}\rangle}}\hskip-2.0pt{\mathinner{\langle{\phi}|}}}{{\mathinner{|{s_{j}}\rangle}}\hskip-2.0pt{\mathinner{\langle{s_{j}}|}}}]=|\mathinner{\langle{\phi|s_{j}}\rangle}|^{2}\leq F$ for each $j$. Therefore,
$|\mathinner{\langle{\psi|s_{j}}\rangle}|^{2}$ $={\rm{Tr~}}[{{\mathinner{|{\psi}\rangle}}\hskip-2.0pt{\mathinner{\langle{\psi}|}}}{{\mathinner{|{s_{j}}\rangle}}\hskip-2.0pt{\mathinner{\langle{s_{j}}|}}}]$ (3.2)
$={\rm{Tr~}}[(P{{\mathinner{|{\psi}\rangle}}\hskip-2.0pt{\mathinner{\langle{\psi}|}}}P){{\mathinner{|{s_{j}}\rangle}}\hskip-2.0pt{\mathinner{\langle{s_{j}}|}}}]\frac{1}{\left\|P\mathinner{|{\phi}\rangle}\right\|_{2}^{2}}$ (3.3)
$={\rm{Tr~}}[{{\mathinner{|{\psi}\rangle}}\hskip-2.0pt{\mathinner{\langle{\psi}|}}}(P{{\mathinner{|{s_{j}}\rangle}}\hskip-2.0pt{\mathinner{\langle{s_{j}}|}}}P)]\frac{1}{|\mathinner{\langle{\psi|\phi}\rangle}|^{2}}$ (3.4)
$={\rm{Tr~}}[{{\mathinner{|{\psi}\rangle}}\hskip-2.0pt{\mathinner{\langle{\psi}|}}}{{\mathinner{|{s_{j}}\rangle}}\hskip-2.0pt{\mathinner{\langle{s_{j}}|}}})]\frac{1}{|\mathinner{\langle{\psi|\phi}\rangle}|^{2}}$ (3.5)
@@
By Theorem 1, we get that $\beta\leq\sqrt{e}\cdot(2k)^{\frac{2k+1}{2}}$. Now, if $\beta=2^{\Omega(n)}$, then $(2k+1)\log(2k)=\Omega(n)\implies k=\Omega\left(\frac{n}{\log n}\right)$
@@ 4 Magic Monotones and Barnes Wall Lattices
[…]
@@
Note that the $2$-norm is equivalent to the norm specified in [kliuchnikov2024stabilizer] on $\mathbb{Q}[i]$. In our results we will restrict to $\mathbb{Q}[i]$.
@@ 4.1 Properties of $\mathcal{N}$ and $\mathcal{N}_{\delta}$
A vector $\mathinner{|{\phi}\rangle}$ is proportionally minimal in $BW_{n}$ if $\mathinner{|{\phi}\rangle}\in BW_{n}$ and for all $0<|c|<1$, $c\mathinner{|{\phi}\rangle}\notin BW_{n}$.
@@
$\forall a,b\in\mathbb{Z}[i]:ac\mathinner{|{\phi}\rangle}+b\mathinner{|{\phi}\rangle}\in BW_{n}$
@@
    $\mathcal{N}$ and $\mathcal{N}_{\delta}$ is invariant under the action of the Clifford group (Also see [kliuchnikov2024stabilizer]).
@@
    $\mathcal{N}(\mathinner{|{\phi}\rangle})\geq 1$, with equality only for stabilizer states. .
@@
    $\mathcal{N}$ is non-increasing under uniform probability $1$-qubit Pauli measurements (post-selecting on any outcome). If it indeed does decrease, it has to divide the original value.
@@
(i) and (ii) follow directly from the invariance of $BW_{n}$ under the action of the Clifford group. [kliuchnikov2024stabilizer].
@@
Now, (iv) follows directly from $(ii)$ and $(iii)$. For (v), we use a technique from [kliuchnikov2024stabilizer]. For any state $\mathinner{|{\phi}\rangle}$, let $\mathinner{|{\phi^{\prime}}\rangle}$ be its proportionally minimal vector. Now, we can write it as, for some vectors $\mathinner{|{v}\rangle},\mathinner{|{w}\rangle}$,
@@
For each such $\alpha$, notice that $(1+i)\mathinner{|{0}\rangle}\mathinner{|{\alpha}\rangle}$ and $(1+i)\mathinner{|{1}\rangle}\mathinner{|{\alpha}\rangle}$ both lie in the dual of $BW_{n}$. Taking their inner products with $\mathinner{|{\phi^{\prime}}\rangle}$, we get that $i\mathinner{\langle{\alpha|v}\rangle}$ and $i\mathinner{\langle{\alpha|w}\rangle}$ respectively, proving that $\mathinner{|{v}\rangle},\mathinner{|{w}\rangle}\in BW_{n-1}$.
@@
Note that a similar property to Theorem 39 $(v)$ holds for the dyadic monotone [dyadic].
@@ 4.2 Connection Between $\mathcal{N}_{\delta}$ and Stabilizer Ranks
[…]
@@
Here the $\{\}$ braces represent the fractional part of $c_{i}$, and $\lfloor{c_{i}}\rfloor$ refers to the greatest integer function. Now, define $k$ pairs of random variables as follows: for each $i\in[k]$, $(A_{i},X_{i})$ is defined as:
@@
[…]
@@ 4.3 Relating $\mathcal{N}$ and $CS$-count for exact state preparation
Figure 1: CSCS-gate injection using the magic state $\mathinner{|{CS}\rangle}$.
@@
$\mathcal{N}(\mathinner{|{\psi}\rangle})\,{\big|}2^{m}$
@@ 5 Fidelity Amplification
$\frac{{\rm{Tr~}}[(I-{{\mathinner{|{\psi}\rangle}}\hskip-2.0pt{\mathinner{\langle{\psi}|}}})\,{{\mathinner{|{\psi^{\prime}}\rangle}}\hskip-2.0pt{\mathinner{\langle{\psi^{\prime}}|}}}]}{{\rm{Tr~}}[{{\mathinner{|{\psi}\rangle}}\hskip-2.0pt{\mathinner{\langle{\psi}|}}}\,{{\mathinner{|{\psi^{\prime}}\rangle}}\hskip-2.0pt{\mathinner{\langle{\psi^{\prime}}|}}}]}$ (5.1)
This can very easily be related to fidelity when $\left\|\mathinner{|{\phi^{\prime}}\rangle}\right\|_{2}=1$, and the expression turns out to be $\frac{1-F(\mathinner{|{\psi}\rangle},\mathinner{|{\phi}\rangle})}{F(\mathinner{|{\psi}\rangle},\mathinner{|{\phi}\rangle})}$.
Observation 45.
@@
$\frac{{\rm{Tr~}}[(I-({{\mathinner{|{H}\rangle}}\hskip-2.0pt{\mathinner{\langle{H}|}}})^{\otimes n})\,{{\mathinner{|{\psi}\rangle}}\hskip-2.0pt{\mathinner{\langle{\psi}|}}}]}{{\rm{Tr~}}[({{\mathinner{|{H}\rangle}}\hskip-2.0pt{\mathinner{\langle{H}|}}})^{\otimes n}\,{{\mathinner{|{\psi}\rangle}}\hskip-2.0pt{\mathinner{\langle{\psi}|}}}]}$ (5.2)
@@
with ${\rm{Tr~}}[{{\mathinner{|{\phi}\rangle}}\hskip-2.0pt{\mathinner{\langle{\phi}|}}}]=\epsilon$, and $\mathinner{\langle{\phi}|}\mathinner{|{H}\rangle}^{\otimes n}=0$.
@@
with $\frac{I+C}{2}\mathinner{|{\phi}\rangle}$ orthogonal to $\mathinner{|{H}\rangle}^{\otimes n}$. Define the $LHS$ above to be $\mathinner{|{\psi^{\prime\prime}}\rangle}$. Note that $\chi(\mathinner{|{\psi^{\prime\prime}}\rangle})\leq 2\chi(\mathinner{|{\psi^{\prime}}\rangle})\leq 2k$. Also, note that ${\rm{Tr~}}[{{\mathinner{|{\psi^{\prime}}\rangle}}\hskip-2.0pt{\mathinner{\langle{\psi^{\prime}}|}}}\,({{\mathinner{|{H}\rangle}}\hskip-2.0pt{\mathinner{\langle{H}|}}})^{\otimes n}]={\rm{Tr~}}[{{\mathinner{|{\psi^{\prime\prime}}\rangle}}\hskip-2.0pt{\mathinner{\langle{\psi^{\prime\prime}}|}}}\,({{\mathinner{|{H}\rangle}}\hskip-2.0pt{\mathinner{\langle{H}|}}})^{\otimes n}]$. Thus, the denominator of Equation 5.2 is the same for both $\mathinner{|{\psi^{\prime}}\rangle}$ and $\mathinner{|{\psi^{\prime\prime}}\rangle}$.
Now, what we expect to happen is that $\frac{I+C}{2}\mathinner{|{\phi}\rangle}$ will have a lot of cancellations if written out in the $\mathinner{|{H}\rangle},\mathinner{|{H^{\perp}}\rangle}$ basis. This is because, for each such basis element, except $\mathinner{|{H}\rangle}^{\otimes n}$ (which by orthogonality has coefficient 0 in the expansion of $\mathinner{|{\phi}\rangle}$), it is an eigenvector of $\frac{I+C}{2}$, with the corresponding eigenvalue being $1$ or 0 with equal probability. We also know that for any vector $\mathinner{|{x}\rangle}$, we have that $\mathinner{\langle{x|x}\rangle}$ is the sum of squares of the coefficients of these basis elements. Thus, we have that for $\mathinner{|{\phi^{\prime}}\rangle}=\frac{I+C}{2}\mathinner{|{\phi}\rangle}$, $\mathbb{E}[\mathinner{\langle{\phi^{\prime}|\phi^{\prime}}\rangle}]=\frac{1}{2}\mathinner{\langle{\phi|\phi}\rangle}=\frac{\epsilon}{2}$, where the randomness is over the choices of $C$. Therefore, there exists a $C$ for which $\mathinner{\langle{\phi^{\prime}|\phi^{\prime}}\rangle}$ is at most $\frac{\epsilon}{2}$. This gives us our result. ∎
@@
Note that if we apply the above result to $\mathinner{|{0}\rangle}^{\otimes n}$, we retrieve back the $O(2^{0.23n})$ rank approximate stabilizer decomposition of $\mathinner{|{H}\rangle}^{\otimes n}$ from [Bravyi_2016]. In fact, our technique can be thought of as generalization for their construction.
5.1 Relation to Lattice Approximations
Here, we discuss the relation of $BW_{n}$ lattice approximation of quantum states, where we say that a lattice vector $v$ on $BW_{n}$ approximates $\mathinner{|{\psi}\rangle}$ if for some $c\in\mathbb{C}^{*}$, $\left\|cv-\mathinner{|{\psi}\rangle}\right\|_{2}$ is small. This is consistent with the notion of approximation used in Definition 35.
In the proof of Lemma 46, we are summing up $I\mathinner{|{\psi^{\prime}}\rangle}$ and $C\mathinner{|{\psi^{\prime}}\rangle}$ where $C$ is a uniformly random $n$-fold tensor product of $H$ and II. [^3] We can ignore the factor of 1/2 as we effectively ignore scalar multiplication while computing the Barnes wall norm $\mathcal{N}$ Now, by Fact 27, we know that whenever $C$ has even number of $H$s, $C$ also maps $BW_{n}$ to itself . Therefore, in such instances, if $\mathinner{|{\psi^{\prime}}\rangle}$ lies on $BW_{n}$, then so does $(I+C)\mathinner{|{\psi^{\prime}}\rangle}$. Therefore, if we restrict to $C$ having an even number of $H$s, the proof of Lemma 46 also acts as a way to jump from one lattice approximation of $\mathinner{|{H}\rangle}^{\otimes n}$ to another.
It is easily seen that if $C$ is chosen uniformly at random from the set of $n$-fold tensor products of $H$ and $I$ with an even number of $I$’s, then the probability of cancelling out components orthogonal to $\mathinner{|{H}\rangle}^{\otimes n}$ in the expansion of $\mathinner{|{\psi^{\prime}}\rangle}$, in the basis of tensor products of $\mathinner{|{H}\rangle},\mathinner{|{H^{\perp}}\rangle}$, is $\frac{1}{2}$ for each of the components except for the component along $\mathinner{|{H^{\perp}}\rangle}^{\otimes n}$. Because of this final caveat, we can not in general say that the new vector on average has half the relative error with respect to $\mathinner{|{H^{\otimes n}}\rangle}$, what we can say is the the relative error is halved with respect to the state $\mathinner{|{\psi^{\prime\prime\prime}}\rangle}$, which is the normalised projection of $\mathinner{|{\psi^{\prime}}\rangle}$ onto the space spanned by $\{\mathinner{|{H}\rangle}^{\otimes n},\mathinner{|{H^{\perp}}\rangle}^{\otimes n}\}$. Also note that $\mathinner{|{\psi^{\prime\prime\prime}}\rangle}$ does not change after sequencial application of this procedure, i.e, the projection of $(I+C)\mathinner{|{\psi^{\prime}}\rangle}$ on to $\mathrm{span}\{\mathinner{|{H}\rangle}^{\otimes n},\mathinner{|{H^{\perp}}\rangle}^{\otimes n}\}$ is also aligned along $\mathinner{|{\psi^{\prime\prime\prime}}\rangle}$. Therefore, for the special case when $\mathinner{|{\psi^{\prime\prime\prime}}\rangle}$ is a very good approximation for $\mathinner{|{H}\rangle}^{\otimes n}$, this modified fidelity amplification procedure still works quite well.
To see why this special case is important, notice that for the setup in Corollary 9, if $\mathinner{|{\psi}\rangle}$ is a good approximation for $\mathinner{|{H}\rangle}^{\otimes m}$, then $\mathinner{|{\psi}\rangle}^{\otimes n}$ will naturally have a very small component along $\mathinner{|{H^{\perp}}\rangle}^{\otimes mn}$, relative to the component along $\mathinner{|{H}\rangle}^{\otimes mn}$.
In particular, the component of the lattice vector $\mathinner{|{0}\rangle}^{\otimes n}$ on to $\mathinner{|{H^{\perp}}\rangle}^{\otimes n}$ is exponentially small relative to it’s component along $\mathinner{|{H}\rangle}^{\otimes n}$. Thus, $O(2^{0.23n})$-rank approximate decomposition for $\mathinner{|{H}\rangle}^{\otimes n}$ obtained from Corolllary 9 or from [Bravyi_2016] can be specialized as a $BW_{n}$ vector approximation for $\mathinner{|{H}\rangle}^{\otimes n}$. Moreover, it can be checked that on average, the square of the length of the lattice vector obtained (i.e, its Barnes wall norm $\mathcal{N}$) scales linearly in the rank. This shows that best known approximations for $\mathinner{|{H}\rangle}^{\otimes n}$ are effectively lattice approximations which asymptotically match the bound in Theorem 7.
Lemma 49 (Informal).
Restricting the randomized $O(2^{0.23n})$-rank decomposition for $\mathinner{|{H}\rangle}^{\otimes n}$ obtained by applying the Fidelility amplification procedure $0.23n$ times to only lattice approximations leads to lattice vector approximations for $\mathinner{|{H}\rangle}^{\otimes n}$ with average squared-length (i.e, the Barnes wall norm $\mathcal{N}$) $O(2^{0.23n})$,
Proof (Sketch).
We split up the problem in to the problem of finding the square of lengths of components along the orthonormal basis formed by tensor products of $\mathinner{|{H}\rangle}$ and $\mathinner{|{H^{\perp}}\rangle}$. This suffices as the square of the lengths of the components sum up to the square of the length of the vector itself.
Suppose $\mathinner{|{\psi}\rangle}$ is the random vector on $BW_{n}$ obtained at the end (after applying 0.23n iterations of the fidelity amplification procedure).
[…]
The length-squared of the component along $\mathinner{|{H^{\perp}}\rangle}^{\otimes n}$ is negligible compared to the above.
For the component orthogonal to the space spanned by $\mathinner{|{H}\rangle}^{\otimes n}$ and $\mathinner{|{H^{\perp}}\rangle}^{\otimes n}$, we note that a basis for this subspace is the set of tensor products of $\mathinner{|{H}\rangle}$ and $\mathinner{|{H^{\perp}}\rangle}$, with $\mathinner{|{H}\rangle}^{\otimes n}$ and $\mathinner{|{H^{\perp}}\rangle}^{\otimes n}$ removed.
For each such basis element, the length squared of the component along $\mathinner{|{\psi}\rangle}$ is quadrupled with probability 1/2, and sent to 0 otherwise. So, on average, the length squared is doubled.
Hence, overall, the length squared component along the space orthogonal to $\mathinner{|{H}\rangle}^{\otimes n}$ and $\mathinner{|{H^{\perp}}\rangle}^{\otimes n}$ is on average $O(2^{0.23n})$ times the initial value , i.e, the length squared of the component of $\mathinner{|{0}\rangle}^{\otimes n}$ on to this space, which is clearly atmost 1. Therefore, the length squared of the lattice vector $\mathinner{|{\psi}\rangle}$ on average is $O(2^{0.23n})$.